I need help solving each system. Thank you.
x-2y-z=3
2x+y-2z=1
2x+3y=3

Answer by lwsshak3(11628)

x-2y-z=3
2x+y-2z=1
2x+3y=3
..
2x-4y-2z=6 (mult 1st eq by 2)
2x+y-2z=1
subtract
-5y=5
y=-1
..
2x+3y=3
2x-3=3(sub (-1) for y)
2x=6
x=3
..
x-2y-z=3
3+2-z=3 (sub (-1) for y and (3) for x)
z=2
Check:
x-2y-z=3+2-2=3
2x+y-2z=4-1-2=1
2x+3y=3=6-3=3
check ok